OpenStudy (anonymous):

How would i show 2^n >n^2 by induction?

OpenStudy (asnaseer):

this doesn't look correct. take the case of n=2.

OpenStudy (anonymous):

but its true after n=5

OpenStudy (asnaseer):

so what does the question state exactly?

OpenStudy (anonymous):

Exactly: 2^n >n^2 if n is element of N and n is large enough

OpenStudy (asnaseer):

ok, so you can show its true for n=5. then assume its true for n=k \((k\ge5)\):\[2^k>k^2\]and finally prove its also true for n=k+1

OpenStudy (anonymous):

yes, im stuck on proving that n=k+1

OpenStudy (asnaseer):

ok, so for n=k+1, we get:\[2^{k+1}=2\times2^k\]and we assumed:\[2^k\gt k^2\]therefore:\[2^{k+1}=2\times2^k\gt2\times k^2\]

OpenStudy (asnaseer):

so you just need to prove that:\[2\times k^2\gt(k+1)^2\]for \(k\ge5\)

OpenStudy (asnaseer):

actually you just need to prove:\[2\times k^2\ge(k+1)^2\]

OpenStudy (asnaseer):

for \(k\ge5\)

OpenStudy (asnaseer):

if you expand \((k+1)^2\) then this reduces to proving:\[k^2\ge2k+1\]for \(k\ge5\)

OpenStudy (anonymous):

In this case, would show that LHS=RHS or LHS>RHS?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

try finding what value of k makes this true:\[k^2-2k-1\ge0\]

OpenStudy (anonymous):

i mean do i show that LHS=RHS or LHS>RHS? do i choose eithe?

OpenStudy (asnaseer):

you need to show BOTH

OpenStudy (anonymous):

how did get \[k ^{2} \ge2k +1?\]

OpenStudy (asnaseer):

\[k^2-2k-1\ge0\]implies:\[(k-1)^2-2\ge0\]\[(k-1)^2\ge2\]therefore:\[k-1\ge\sqrt{2}\]therefore:\[k\ge1+\sqrt{2}\]

OpenStudy (asnaseer):

we has to prove that:\[2\times k^2\ge(k+1)^2\]therefore:\[k^2+k^2\ge k^2+2k+1\]one of the \(k^2\) terms cancels out

OpenStudy (anonymous):

I am assuming \[RHS=k^2 +2k+1\]

OpenStudy (anonymous):

and from that i use the assumption

OpenStudy (asnaseer):

?

OpenStudy (anonymous):

When i prove by induction, i prove the LHS and RHS seperately...

OpenStudy (asnaseer):

let me try and summarise the steps for you...

OpenStudy (asnaseer):

we first showed that:\[2^5>5^2\]which shows that:\[2^n\gt n^2\]for \(n=5\) - agreed?

OpenStudy (anonymous):

Like, i want to know how from \[RHS=k^2+2k+1 \] would i incorporate the assumption

OpenStudy (anonymous):

yes i agree

OpenStudy (asnaseer):

ok, so next step was to /assume/ this relation is true for some \(n=k\) where \(k\ge5\), so we get:\[2^k\gt k^2\]agreed?

OpenStudy (anonymous):

yes i agree

OpenStudy (asnaseer):

ok, so next step was to then prove, assuming its true for \(n=k\), that it's also true for \(n=k+1\) where \(k\ge5\). so all we need to prove is:\[2^{k+1}\gt (k+1)^2\]for \(k\ge5\) - agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so if we look at just the term \(2^{k+1}\) we see that this can be written as:\[2^{k+1}=2\times2^k\]but, we have already assumed that:\[2^k\gt k^2\]therefore we can assert that:\[2^{k+1}=2\times2^k\gt2\times k^2\]agreed?

OpenStudy (anonymous):

so you are saying that \[RHS=2^k \times 2 and thus RHS>2timesk^2\]

OpenStudy (asnaseer):

you know that:\[2^{k+1}=2\times2^k\]correct?

OpenStudy (anonymous):

so you just replaced 2^k with k^2 and the = must now be > after substituting?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

yes, i agree

OpenStudy (asnaseer):

ok, so the last step we got to was that we have so far proved that:\[2^{k+1}\gt2\times k^2\]

OpenStudy (asnaseer):

now, what we /need/ to prove is that:\[2^{k+1}\gt(k+1)^2\]correct?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

ok, so if we can prove that:\[2\times k^2\ge(k+1)^2\]then we have proved that:\[2^{k+1}\gt(k+1)^2\]agreed?

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

good - so lets see what values of k satisfy the equation:\[2\times k^2\ge(k+1)^2\]

OpenStudy (asnaseer):

we get:\[k^2+k^2\ge(k+1)^2\]\[\cancel{k^2}+k^2\ge \cancel{k^2}+2k+1\]\[k^2\ge 2k+1\]\[k^2-2k-1\ge0\]\[(k-1)^2-2\ge0\]\[(k-1)^2\ge2\]\[k-1\ge\sqrt{2}\]\[k\ge1+\sqrt{2}\]

OpenStudy (asnaseer):

but we wanted this to be true for \(k\ge5\) so we have now proved what we wanted.

OpenStudy (anonymous):

\[2k <2^{k}\] if k<0 conjecture is true as LHS <0 andRHS>0 For k=1 , 2*1-2^1 is true assume true for k=n \[2n<^{2^{n}}\] then \[2n \times2<2^{n}\times2\] \[4n <2^{n+1}\] \[2n+2n <2^{n+1}\] so \[2n+2<2^{n+1}\] as \[2n >2 for n >1\] i.e. \[2(k+1)<2^{k+1}\] if conjecture is true fork=n then it is true for k=n+1 proved

OpenStudy (anonymous):

you mean the fact that 5 is greater than 1+ sqrt(2) is proven?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

and that is all?

OpenStudy (asnaseer):

yes :)

OpenStudy (anonymous):

yu guys made it very complicated :/

OpenStudy (anonymous):

hang on, i agree with you now :) but what i mean is, have we shown that LHS=RHS?

OpenStudy (asnaseer):

@Tushara we were not trying to prove \(2k <2^{k}\)

OpenStudy (anonymous):

oh oops sowwie

OpenStudy (asnaseer):

we have proved:\[2^n\gt n^2\]for \(n\ge5\)

OpenStudy (anonymous):

ok sweet

OpenStudy (asnaseer):

@Tushara np - its good to have other people double checking your work - we can all make mistakes from time to time :)

OpenStudy (anonymous):

like what i am doing: